OpenStudy (anonymous):

What is the slope 2x+6

jimthompson5910 (jim_thompson5910):

y = 2x+6 is in the form y = mx+b where m = 2 Since the slope of y = mx+b is 'm' and m = 2, the slope of y = 2x+6 is 2

OpenStudy (mertsj):